Since the birth of their daughter, the Tranters have deposited $450 every 3 months on an education savings plan. The interest rate is 7.5%/a compounded quarterly. What is the plans value when their daughter turns 17?
Answer by jim_thompson5910(35256)

Use this formula:

FV = P*( (1+i)^(n) - 1 )/i

where,

P = amount deposited per period
i = interest rate per period
n = number of periods


In this case,

P = 450
i = 0.075/4 = 0.01875
n = 4*17 = 68

------------------------------------------------------------------

FV = P*( (1+i)^(n) - 1 )/i

FV = 450*( (1+0.01875)^(68) - 1 )/0.01875

FV = 60880.8110324299

FV = 60880.81


They will have $60,880.81 in the account 17 years later (assuming they keep up with every deposit)


See this article for more details:
http://www.investopedia.com/terms/f/future-value-annuity.asp